Consider A, B, and C three points of the plan. Show that det(AB, AC) = det(BC, BA) = det (CA, CB)

a) Using a geometrical approach

b) Using the property of antisymmetry of the determinant



1
Expert's answer
2022-01-24T15:41:04-0500

a)

The magnitude of the cross product can be interpreted as the positive area of the parallelogram having a and b as sides.

in our case:

"det(AB, AC) = det(BC, BA) = det (CA, CB)=2(area\ of\ \Delta ABC)"


b)

antisymmetry property of determinants states that interchanging two rows of a determinant changes the sign of the determinant


then:


"det(AB, AC)=det(AB, AB+BC)=det(AB, AB)+det(AB,BC)="

"=det(AB,BC)"


"det(BC, BA)=-det(BC, AB)=det(AB,BC)"


"det (CA, CB)=-det (CA, BC)=-det (-(AB+BC), BC)="

"=det (AB+BC, BC)=det(AB,BC)+det(BC,BC)=det(AB,BC)"


so, "det(AB, AC) = det(BC, BA) = det (CA, CB)"